y=5x+7,y- x=15


1
Expert's answer
2019-10-16T12:17:08-0400

"\begin{cases}
 y=5x+7 \\
y-x=15 
\end{cases}"

Express "y" in terms of "x" using the first equation and substitute for "y" into the second equation to find values of "x"

and we have:


"5x+7-x=15\\
5x-x=15-7\\
4x=8\\
x=2"

after that we substitute the computed value of "x" into the first equation to find the value of "y"


"y=5*2+7\\
y=17"

Answer: "x=2" ; "y=17"